An electrical firm manufactures light bulbs that have a length life of that is approximately normally distributed with a mean of 800 hours and a standard deviation of 40 hours. Test the hypothesis that 𝜇 = 800 ℎ𝑜𝑢𝑟𝑠 against the alternative μ ≠ 800 hours if random sample of 30 bulbs has an average life of 788 hours. Use a 0.01


1
Expert's answer
2022-01-18T13:46:41-0500

"H_0: \mu = 800 \\

H_1: \mu ≠ 800 \\

n = 30 \\

\sigma = 40 \\

\bar{x} = 788 \\

α = 0.01"

Test-statistic

"Z = \frac{\bar{x} - \mu}{\sigma / \sqrt{n}} \\

Z = \frac{788 -800}{40 / \sqrt{30}} = -1.643"

Two-tailed test.

Reject H0 is Z ≤ -2.575 or Z≥ 2.575.

Since Z = -1.643 exceeds -2.575, reject the null hypothesis at the 1 % significance level.

We can conclude that "\mu ≠ 800" .